If a black body radiates 20 calories per second at 227°C, it will radiate at 727°C

Grade:12th pass

1 Answers

Arun
6 years ago
Energy emitted by a body is directly proportional to T^4 (Stefan's law) 
so E1/E2 = (T1/T2)^4 
x/20 = (727+273/227+273)^4 ..Absolute temp in Kelvin 
x = 16*20 = 320 calories per second